TPC weekly

2021-02-24 15:00
America/New York
Wednesday, 24 February 2021
, at 20:00 (GMT), duration : 00:00

- the west laser had a LV power supply with a bad fuse. Alexie got an access fand fixed in < 1 hour
- Alexei monitor gas and will adjust methan concentration slightly.
- new hot channel in sector1. Have asked Tonko to mask out

FF request.
Gene summarized the FF request as it had emerged from the calibration review. It is important to run FF setting to seperate alignment, E*B distortins
but having data from both polarities. Yuri & Gene confirm that the 20M RF is a good sample from RFF and it was concluded we should request a change as soon as practical and on Wednesday access day at latest. Alexei brought up request to at at the time of switch that a high stat laser run ~20k is taken jest before magnet change and another one is taken with FF after the change, and magnet stabilized.

Yuri showed the dedx for his run 20 calibration for collider data and fixed target.There is some room for improvement in particular for the FXT target data.
Yuri showed that the readout of the magnets current fluctuate on a 15 sec time sacle between two distinct values. Current code only has one magnet value per run so it gives rise to distinct bands of say K0_as there is a scale factor between current read and Field suggest to use average value for future.

Gene discussed a Minor problem in sector alignment distortion correction as shown in attached slide on web page.
Looked at run 18 distortion maps and notice a missing featue that he expected namely spikes missing between inner and outer sector. This is due to an un intentional change of code in 2014. Demonstrated with differences that these effect are likely small. One measure of #primary tracks is sub-promille.

15:00Run XX dE/dx calibration ( 00:05 ) 0 filesYuri Fisyak
15:05Run XXI Ks0 mass and STAR main magnet current ( 00:05 ) 0 filesYuri Fisyak
15:10Minor problem in sector alignment distortion correction ( 00:20 ) 0 filesG. Van Buren (BNL)